Suppose peas are planted in Indiana on May 1. If the peas need 1200 growing degree days before they can be picked, and if the average maximum temperature for May and June is 80°F and the average minimum is 60°F, on about what date will the peas be ready to pick? (Assume a base temperature of 40°F.)

a. May 30

b. June 9

c. June 29
Answer #1

average mean temperature for may and june = (80+60)/2 = 70

given base temperature of 40

difference = 70-40 =30

and 1200 growing degree days needed

so total days will be = 1200/30 = 40 days

which means june 9 is correct answer